# FDF: Flexible Decoupled Framework for Time Series Forecasting with Conditional Denoising and Polynomial Modeling

### v1

@@ Abstract.
Time series forecasting is vital in numerous web applications, influencing critical decision-making across industries. While diffusion models have recently gained increasing popularity for this task, we argue they suffer from a significant drawback: indiscriminate noise addition to the original time series followed by denoising, which can obscure underlying dynamic evolving trend and complicate forecasting. To address this limitation, we propose a novel flexible decoupled framework (FDF) that learns high-quality time series representations for enhanced forecasting performance. A key characteristic of our approach leverages the inherent inductive bias of time series data by decomposing it into trend and seasonal components, each modeled separately to enable decoupled analysis and modeling. Specifically, we propose an innovative Conditional Denoising Seasonal Module (CDSM) within the diffusion model, which leverages statistical information from the historical window to conditionally model the complex seasonal component. Notably, we incorporate a Polynomial Trend Module (PTM) to effectively capture the smooth trend component, thereby enhancing the model’s ability to represent temporal dependencies. Extensive experiments validate the effectiveness of our framework, demonstrating superior performance over existing methods and highlighting its flexibility in time series forecasting. We hope our work can bring a new perspective for time series forecasting. We intend to make our code publicly available as open-source in the future.
Diffusion Model; Time Series Forecasting; Decoupled modeling
@@ 1. INTRODUCTION
Time series forecasting is a fundamental task in various domains, playing a pivotal role in driving informed decision-making across industries (Kashpruk et al., 2023). It involves analyzing data captured in a historical window to model the dynamic evolution of underlying processes and predict future trends over a predicted window (Montgomery et al., 2015). In web applications, accurate time series forecasting is essential for tasks such as demand prediction in e-commerce platforms (Trastour et al., 2002), load balancing in cloud computing services (Kilcioglu et al., 2017), and user engagement forecasting (Drutsa et al., 2015) in social media analytics.
Over the years, a multitude of approaches have been proposed to tackle this problem, ranging from classical statistical models like Autoregressive Integrated Moving Average (ARIMA) (Asteriou and Hall, 2011) to advanced deep learning techniques employing recurrent neural networks (RNNs)(Lin et al., 2023; Lai et al., 2018), convolutional neural networks (CNNs)(Bai et al., 2018; Liu et al., 2022), and transformer-based networks (Zhou et al., 2021; Wu et al., 2021). Despite their effectiveness, these previous methods often lack consideration for the inherent complexities of time series data, such as non-linear temporal dependencies, continuous value ranges, and the probabilistic nature of future events (Kolambe and Arora, 2024). They may struggle to capture the uncertainty associated with future forecasts or to model complex patterns like heteroscedasticity and non-stationarity. Probabilistic models have emerged as promising alternatives, offering a principled way to quantify uncertainty and model intricate data distributions.
@@
To overcome these challenges, we propose a novel flexible decoupled framework(FDF) for time series forecasting that leverages the inductive bias of time series decomposition. By separating the time series into trend and seasonal components, we enable decoupled modeling tailored to each component’s characteristics. Specifically, we employ meticulously designed conditional denoising seasonal module (CSDM) to capture the complex seasonal component, introducing an innovative denoising strategy conditioned on statistical information derived from the historical window. Concurrently, we model the smooth trend component using polynomial trend module (PTM), which are adept at capturing gradual changes over time. This decoupling method enhances the model’s ability to capture temporal dependencies and complex patterns without sacrificing critical trend information.
The key contributions of this work are:
@@ 2.1. Time Series Forecasting
Time series forecasting has evolved from traditional statistical methods to advanced deep learning approaches for more accurate predictions. Early models like VAR (Watson, 1994) and ARIMA (Asteriou and Hall, 2011) were widely used for linear temporal dependencies, but their limitations in capturing complex nonlinear patterns prompted a shift toward deep learning. Recurrent neural networks (RNNs), such as LSTNet (Lai et al., 2018) and SegRNN (Lin et al., 2023), have become popular for modeling sequential dependencies. Concurrently, convolutional neural networks (CNNs) like TCN (Bai et al., 2018) and SCINet (Liu et al., 2022) excel at capturing both local and global temporal features through convolutional filters. In addition, graph-based models such as MTGNN (Wu et al., 2020) and FourierGNN (Yi et al., 2024) have improved forecasting by modeling interdependencies between time series. More recently, Transformer-based models, including Informer (Zhou et al., 2021) and Autoformer (Wu et al., 2021), have advanced the field by utilizing self-attention mechanisms to efficiently capture long-range dependencies, proving particularly effective in high-dimensional and long-horizon tasks. Interestingly, simpler linear models like DLinear (Zeng et al., 2023) and TSMixer (Ekambaram et al., 2023) have shown that, by focusing on long-term trends, well-tuned linear layers can outperform more complex architectures, emphasizing efficiency and scalability. While current deep learning methods have strong memory capabilities, they often struggle with modeling complex and volatile fluctuations, presenting challenges in capturing nuanced time series dynamics.
@@ 2.2. Diffusion Models
Figure 2. The Flexible Decoupled Framework (FDF) improves forecasting accuracy by decoupling. The Polynomial Trend Module (PTM) models the trend component to capture different patterns, while the Conditional Denoising Seasonal Module (CDSM) refines the seasonal component through a diffusion process conditioned on historical data.
@@ 3.6.2. Optimization
Table 1. Overview of the dataset characteristics, including dimension, sampling frequency, and the target series $T$ used for prediction in the experiments.
@@
4. EXPERIMENTS
In this section, we assess the effectiveness of the proposed FDF framework through comprehensive experimental comparisons.
Table 2. Time series forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined. CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
@@ 4.2.1. Main results
We assess the performance of the Flexible Decoupled Framework (FDF) across several real-world datasets, as summarized in Table 2. FDF consistently outperforms state-of-the-art models, demonstrating notable improvements in Mean Squared Error (MSE) and Mean Absolute Error (MAE) over methods such as Diffusion-TS, D3VAE, and PatchMixer, confirming its effectiveness for time series forecasting. On challenging datasets like ETTh and ETTm, FDF adeptly manages varying time resolutions and intensities. In the Weather dataset, FDF sustains high accuracy across extended forecasting windows, underscoring its strength in forecasting tasks. While the Electricity dataset presents challenges due to its higher dimensionality, FDF still maintains competitive performance. The Conditional Denoising Seasonal Module (CDSM) captures complex fluctuations, while the Polynomial Trend Module (PTM) effectively preserves trends, capturing both linear and nonlinear dynamics. Together, these modules enhance FDF’s adaptability across diverse forecasting scenarios, a capability that is also validated by the more detailed experimental results presented in Appendix C.1.
@@ 4.2.3. Decoupling Variants Experiments
  • FDFa: This variant bypasses the decoupling process, directly applying the CDSM to the original time series data.
  • FDFb: In this variant, the roles of PTM and CDSM are reversed, with PTM predicting the seasonal component and CDSM predicting the trend component.
We evaluate the impact of the decoupling strategy on the performance of our proposed FDF model using the ETTh2 and ETTm2 datasets. As shown in Figure 5, we compare the following two model variants:
  • FDFa: This variant bypasses the decoupling process, directly applying the CDSM to the original time series data.
  • FDFb: In this variant, the roles of PTM and CDSM are reversed, with PTM predicting the seasonal component and CDSM predicting the trend component.
The results demonstrate that the decoupling strategy significantly enhances model performance, with FDF outperforming both ablated models across the datasets. The improvement is more pronounced on ETTh2, particularly in MSE, while gains on ETTm2 are more moderate. This difference can be attributed to the varying sampling granularity between the two datasets. ETTh2 presents clearer and more distinct trends and seasonal patterns, allowing the decoupling strategy to more effectively separate and model these components. In contrast, ETTm2, with its finer sampling granularity, exhibits more complex and overlapping patterns, making it more difficult to isolate trends from seasonal fluctuations, thus leading to smaller performance gains. Despite this, FDF still consistently outperforms the ablated models on both datasets, highlighting its adaptability. FDFa underperforms due to its inability to separate components, while FDFb offers slight improvements but still lags behind FDF, further emphasizing the importance of assigning each module to its respective component. Overall, these findings highlight the crucial role of tailoring each module to capture the distinct characteristics of trends and seasonal patterns, as further elaborated in Appendix C.2. This is especially important for datasets with varying granularities, where such customization is key to achieving optimal forecasting performance.
Figure 6. Parameter sensitivity analysis of FDF.
@@ 4.2.4. Ablation Experiments
FDF1 0.3404 0.3781 0.2320 0.3023
FDF2 0.3433 0.3826 0.2450 0.3121
FDF3 0.3853 0.4215 0.2459 0.3272
FDF4 0.3442 0.3807 0.2335 0.3015
FDF5 0.2300 0.3006 0.2300 0.3006
@@
  • FDF1: Removes the conditional information.
  • FDF2: Replaces the PTM with a simple linear layer.
  • FDF3: Removes instance normalization during both the training and inference phases.
  • FDF4: Removes the conditional information and replaces PTM with a linear layer.
  • FDF5: Removes instance normalization alignment.
@@ 4.2.5. Hyper-parameter sensitiviy analysis
We conduct a hyperparameter sensitivity analysis on the ETTh1 dataset to explore how key design choices in FDF—hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon—affect its performance. As shown in Figure 6, increasing $d_{\text{model}}$ initially enhances performance, reaching an optimal value at $d_{\text{model}}=256$. However, larger values, such as $d_{\text{model}}=512$, result in diminishing returns and potential overfitting, underscoring the importance of balancing model capacity with computational efficiency. Similarly, patch length $H$ significantly impacts model effectiveness. With $H=8$, FDF effectively captures both local dependencies and broader temporal patterns through the CDSM. Larger $H$ values introduce redundancy and decrease efficiency, indicating that careful selection of $H$ is crucial for accurate multi-scale modeling. Finally, Figure 6(c) shows that FDF maintains consistent performance across different prediction horizons. The PTM captures various trends, while the diffusion-based seasonal component adapts to fluctuations, ensuring reliable forecasting across timeframes. Overall, the analysis confirms FDF’s balance between complexity and scalability, validating its suitability for a wide range of forecasting applications.
@@ 4.2.6. PTM & CDSM Joint Modeling Capability Visualization
Figure 7 illustrates the combined performance of the PTM and CDSM sub-modules within the FDF framework. By comparing predicted and actual trend and seasonal components, the framework’s interpretability and modeling capabilities are validated. The FDF framework effectively decouples multidimensional time series, capturing diverse trends and cyclical fluctuations, demonstrating its ability to model complex temporal patterns. In Figure 7 (a), focusing on the Electricity dataset, which exhibits clear periodic and trend characteristics, the FDF framework shows strong performance in trend modeling. PTM accurately captures smooth and stable trends, with predictions closely aligning with the actual trends. This is attributed to PTM’s ability to effectively model the overall temporal structure, allowing it to identify key trend features. Figure 7 (b) highlights CDSM’s ability to capture complex seasonal variations. The predicted seasonal components correspond well to actual patterns, and CDSM compensates for deviations in PTM’s trend predictions, further improving overall accuracy. This demonstrates CDSM’s flexibility in modeling multi-level temporal features, particularly in noisy data with intricate cyclical patterns. In summary, the joint modeling of PTM and CDSM within the FDF framework shows strong complementarity. The modular design allows FDF to handle datasets containing both trend and cyclical components, offering enhanced interpretability and predictive accuracy.
@@ A.2. Datasets
We perform extensive experiments on several real-world time series datasets across diverse domains to assess the performance of our proposed framework. The datasets used in this study are: (i) ETTh and ETTm(Zhou et al., 2021), comprising two years of electricity transformer temperature data from China, recorded at 1-hour and 15-minute intervals, respectively. These datasets offer valuable insights into various trends and complex variations in electricity usage. (ii) Exchange, which includes daily exchange rates from eight countries: Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore. This dataset presents challenges due to the complex dependencies and volatility inherent to financial time series. (iii) Weather [^3] https://www.bgc-jena.mpg.de/wetter/, containing 21 meteorological indicators recorded every 10 minutes from 2020 to 2021, providing data that captures seasonal patterns and sudden weather changes. (iv) Electricity [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, which provides hourly electricity consumption data from 321 clients over two years. This dataset is characterized by complex behaviors influenced by external factors. (v) Wind(Li et al., 2022), containing wind power generation data recorded at 15-minute intervals from 2020 to 2021, reflecting dynamics shaped by daily routines and unforeseen events.
@@ C.1. Main Results
Table 6. MSE and MAE Performance Across Datasets. The best result is in bold, and the second-best is underlined. CSDI run out of memory on Electricity and Diffusion-TS, while D3VAE run out of memory on Electricity.
@@ C.2. Decoupling Variants Experiments
  • FDFa: This variant bypasses the decoupling process, directly applying the CDSM to the original time series data.
  • FDFb: In this variant, the roles of PTM and CDSM are reversed, with PTM predicting the seasonal component and CDSM predicting the trend component.
The decoupling experiments provide key insights into the impact of separating trend and seasonal components on forecasting accuracy, as shown in Table 7. The fully decoupled model consistently outperforms ablated versions, especially on complex datasets like ETTh2, where significant gains in MSE and MAE are observed. For example, the decoupled model achieves an MSE of 0.3393, compared to 0.3839 for the non-decoupled variant (FDFa), highlighting the difficulty of modeling trend and seasonal components together. The role-swapping variant (FDFb), where the Polynomial Trend Module (PTM) and Conditional Denoising Seasonal Module (CDSM) are misaligned, shows only marginal improvements over FDFa but still underperforms the fully decoupled model. For instance, on ETTm2, the MSE improves slightly from 0.2488 to 0.2553, underscoring the importance of correct module assignment. PTM is better suited for capturing smooth trends, while CDSM excels at modeling complex seasonal patterns, and misalignment compromises their effectiveness. These findings reinforce the importance of the decoupling strategy in time series forecasting. Proper separation of trend and seasonal components, combined with correct module alignment, allows each module to focus on its respective temporal patterns, enhancing forecasting accuracy and interpretability. The results confirm that decoupling trend and seasonal patterns is crucial for improved model performance, as validated by the experiments.

### v2

@@ Abstract.
Time series forecasting is vital in numerous web applications, influencing critical decision-making across industries. While diffusion models have recently gained increasing popularity for this task, we argue they suffer from a significant drawback: indiscriminate noise addition to the original time series followed by denoising, which can obscure underlying dynamic evolving trend and complicate forecasting. To address this limitation, we propose a novel flexible decoupled framework (FDF) that learns high-quality time series representations for enhanced forecasting performance. A key characteristic of our approach leverages the inherent inductive bias of time series data of its decomposed trend and seasonal components, each modeled separately to enable decoupled analysis and modeling. Specifically, we propose an innovative Conditional Denoising Seasonal Module (CDSM) within the diffusion model, which leverages statistical information from the historical window to conditionally model the complex seasonal component. Notably, we incorporate a Polynomial Trend Module (PTM) to effectively capture the smooth trend component, thereby enhancing the model’s ability to represent temporal dependencies. Extensive experiments validate the effectiveness of our framework, demonstrating superior performance over existing methods and highlighting its flexibility in time series forecasting. The source code is available at https://github.com/zjt-gpu/FDF.
Time Series Forecasting; Diffusion Model; Decoupled Modeling
@@ 1. INTRODUCTION
Time series forecasting is a fundamental task in various domains, playing a pivotal role in driving informed decision-making across industries (Kashpruk et al., 2023; Liu et al., 2024). It involves analyzing data captured in a historical window to model the dynamic evolution of underlying processes and predict future trends over a predicted window (Montgomery et al., 2015; Cheng et al., 2024). In web applications, accurate time series forecasting is essential for tasks such as demand prediction in e-commerce platforms (Trastour et al., 2002), load balancing in cloud computing services (Kilcioglu et al., 2017), and user engagement forecasting (Drutsa et al., 2015) in social media analytics.
Over the years, a multitude of approaches have been proposed to tackle this problem, ranging from classical statistical models like Autoregressive Integrated Moving Average (ARIMA) (Asteriou and Hall, 2011) to advanced deep learning techniques employing recurrent neural networks (RNNs)(Lin et al., 2023; Lai et al., 2018), convolutional neural networks (CNNs)(Bai et al., 2018; Liu et al., 2022; Cheng et al., 2024), and transformer-based networks (Zhou et al., 2021; Wu et al., 2021; Cheng et al., 2023a). Despite their effectiveness, these previous methods often lack consideration for the inherent complexities of time series data, such as non-linear temporal dependencies, continuous value ranges, and the probabilistic nature of future events (Kolambe and Arora, 2024). They may struggle to capture the uncertainty associated with future forecasts or to model complex patterns like heteroscedasticity and non-stationarity. Probabilistic models have emerged as promising alternatives, offering a principled way to quantify uncertainty and model intricate data distributions.
@@
To overcome these challenges, we propose a novel flexible decoupled framework(FDF) for time series forecasting that leverages the inductive bias of time series decomposition. By separating the time series into trend and seasonal components, we enable decoupled modeling tailored to each component’s characteristics. Specifically, we employ meticulously designed conditional denoising seasonal module (CSDM) to capture the complex seasonal component, introducing an innovative denoising strategy conditioned on statistical information derived from the historical window. Concurrently, we model the smooth trend component using polynomial trend module (PTM), which are adept at capturing gradual changes over time. This decoupling method enhances the model’s ability to capture temporal dependencies and complex patterns without sacrificing critical trend information. The key contributions of this work are:
@@ 2.1. Time Series Forecasting
Time series forecasting has evolved from traditional statistical methods to advanced deep learning approaches for more accurate predictions (Cheng et al., 2023b). Early models like VAR (Watson, 1994) and ARIMA (Asteriou and Hall, 2011) were widely used for linear temporal dependencies, but their limitations in capturing complex nonlinear patterns prompted a shift toward deep learning. Recurrent neural networks (RNNs), such as LSTNet (Lai et al., 2018) and SegRNN (Lin et al., 2023), have become popular for modeling sequential dependencies. Concurrently, convolutional neural networks (CNNs) like TCN (Bai et al., 2018) and SCINet (Liu et al., 2022) excel at capturing both local and global temporal features through convolutional filters. In addition, graph-based models such as MTGNN (Wu et al., 2020) and FourierGNN (Yi et al., 2024) have improved forecasting by modeling interdependencies between time series. More recently, Transformer-based models, including Informer (Zhou et al., 2021) and Autoformer (Wu et al., 2021), have advanced the field by utilizing self-attention mechanisms to efficiently capture long-range dependencies, proving particularly effective in high-dimensional and long-horizon tasks. Interestingly, simpler linear models like DLinear (Zeng et al., 2023) and TSMixer (Ekambaram et al., 2023) have shown that, by focusing on long-term trends, well-tuned linear layers can outperform more complex architectures, emphasizing efficiency and scalability. While current deep learning methods have strong memory capabilities, they often struggle with modeling complex and volatile fluctuations, presenting challenges in capturing nuanced time series dynamics.
@@ 2.2. Diffusion Models
Figure 2. Illustration of the proposed Flexible Decoupled Framework (FDF). The Polynomial Trend Module (PTM) models the trend component to capture different patterns, while the Conditional Denoising Seasonal Module (CDSM) refines the seasonal component through a diffusion process conditioned on historical data.
@@ 3.5. Polynomial Trend Module
3.6. Fusion and Optimization
3.6.1. Fusion
3.6.2. Optimization
4. EXPERIMENTS
Table 1. Overview of the dataset characteristics. Here, $T$ denotes the target series used for prediction.
@@ 4. EXPERIMENTS
Table 2. The forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined. CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
@@ 4.2.1. Main results
We assess the performance of the Flexible Decoupled Framework (FDF) across several real-world datasets, as summarized in Table 2. FDF consistently outperforms state-of-the-art models, demonstrating notable improvements in Mean Squared Error (MSE) and Mean Absolute Error (MAE) over methods such as Diffusion-TS, D3VAE, and PatchMixer, confirming its effectiveness for time series forecasting. On challenging datasets like ETTh and ETTm, FDF adeptly manages varying time resolutions and intensities. In the Weather dataset, FDF sustains high accuracy across extended forecasting windows, underscoring its strength in forecasting tasks. While the Electricity dataset presents challenges due to its higher dimensionality, FDF still maintains competitive performance. The Conditional Denoising Seasonal Module (CDSM) captures complex fluctuations, while the Polynomial Trend Module (PTM) effectively preserves trends, capturing both linear and nonlinear dynamics. Together, these modules enhance FDF’s adaptability across diverse scenarios, a capability that is also validated by the more detailed experimental results in Appendix C.1.
@@ 4.2.3. Decoupling Variants Experiments
  • FDFw/o ${}_{\text{w/o decoupling}}$: This variant bypasses the decoupling process, directly applying the CDSM to the original time series.
[…]
@@ 4.2.4. Ablation Experiments
FDFw/o ${}_{\text{w/o CR}}$ 0.3404 0.3781 0.2320 0.3023
[…]
FDFw/o ${}_{\text{w/o INR}}$ 0.3853 0.4215 0.2459 0.3272
FDFw/o ${}_{\text{w/o CR+PLR}}$ 0.3442 0.3807 0.2335 0.3015
FDFw/o ${}_{\text{w/o INA}}$ 0.2300 0.3006 0.2300 0.3006
@@
  • FDFw/o ${}_{\text{w/o CR}}$: Removes the conditional information.
[…]
  • FDFw/o ${}_{\text{w/o INR}}$: Removes instance normalization during both the training and inference phases.
  • FDFw/o ${}_{\text{w/o CR+PLR}}$: Removes the conditional information and replaces PTM with a linear layer.
  • FDFw/o ${}_{\text{w/o INA}}$: Removes instance normalization alignment.
@@ 4.2.5. Hyper-parameter sensitiviy analysis
We conducted a hyperparameter sensitivity analysis on the ETTh1 dataset to explore the impact of key design choices in FDF—hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon—on its performance. The results show that increasing $d_{\text{model}}$ initially improves performance, peaking at $d_{\text{model}}=256$, but larger values such as $d_{\text{model}}=512$ lead to diminishing returns and potential overfitting, indicating the need to balance model capacity and computational efficiency. Similarly, $H$ significantly affects performance, with $H=8$ effectively capturing both local dependencies and global temporal patterns, while larger values introduce redundancy and reduce efficiency. Finally, FDF maintains consistent performance across different prediction horizons, with the PTM capturing trends and the diffusion-based seasonal component adapting to fluctuations, ensuring reliable forecasting. Overall, the analysis confirms FDF’s balance between complexity and scalability, making it suitable for a wide range of forecasting tasks.
@@ 4.2.6. PTM & CDSM Joint Modeling Capability Visualization
Figure 7 shows the combined performance of the PTM and CDSM sub-modules in the FDF framework. By comparing predicted and actual trend and seasonal components, the framework’s interpretability and modeling capabilities are validated. The FDF effectively decouples multidimensional time series, capturing diverse trends and cyclic patterns, demonstrating its strength in modeling complex temporal dynamics. In Figure 7 (a), applied to the Electricity dataset with clear periodic and trend characteristics, PTM captures smooth, stable trends with predictions closely matching actual trends, thanks to its ability to model the overall temporal structure. Figure 7 (b) showcases CDSM’s ability to capture complex seasonal variations. The predicted seasonal components align well with actual patterns, compensating for PTM’s trend deviations and further improving accuracy. This highlights CDSM’s flexibility in handling noisy data with intricate cyclic features. Overall, the joint modeling of PTM and CDSM within the FDF framework demonstrates strong complementarity, enabling enhanced interpretability and predictive accuracy in datasets with both trend and seasonal components.
@@ A.2. Datasets
We perform extensive experiments on several real-world time series datasets across diverse domains to assess the performance of our proposed framework. The datasets used in this study are: (i) ETTh and ETTm (Zhou et al., 2021), comprising two years of electricity transformer temperature data from China, recorded at 1-hour and 15-minute intervals, respectively. These datasets offer valuable insights into various trends and complex variations in electricity usage. (ii) Exchange (Lai et al., 2018), which includes daily exchange rates from eight countries: Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore. This dataset presents challenges due to the complex dependencies and volatility inherent to financial time series. (iii) Weather [^3] https://www.bgc-jena.mpg.de/wetter/, containing 21 meteorological indicators recorded every 10 minutes from 2020 to 2021, providing data that captures seasonal patterns and sudden weather changes. (iv) Electricity [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, which provides hourly electricity consumption data from 321 clients over two years. This dataset is characterized by complex behaviors influenced by external factors. (v) Wind (Li et al., 2022), containing wind power generation data recorded at 15-minute intervals from 2020 to 2021, reflecting dynamics shaped by daily routines and unforeseen events.
@@ C.1. Main Results
Table 6. The forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined. CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
@@ C.2. Decoupling Variants Experiments
  • FDFw/o ${}_{\text{w/o decoupling}}$: This variant bypasses the decoupling process, directly applying the CDSM to the original time series data.
[…]